OpenStudy (yourboydusty):

Which of the following points are on the line given by the equation y = -3x +5? Check all that apply. A. (-3, 14) B. (3, -14) C. (2, -1) D. (-2, 11)

OpenStudy (3mar):

Well, I am here.

OpenStudy (phi):

do you know what -3*-3 + 5 is ? (use order of operations: first do the multiply)

OpenStudy (yourboydusty):

-4

OpenStudy (phi):

a minus times a minus is a plus (one of those mysteries) -3 * -3 is not -9 try again

OpenStudy (yourboydusty):

9

OpenStudy (phi):

ok so what is -3*-3 + 5 ?

OpenStudy (yourboydusty):

14

OpenStudy (3mar):

Correct

OpenStudy (yourboydusty):

So the answer would be A

OpenStudy (phi):

what you just did was test (-3, 14) (choice A) we put in -3 for x in -3*x + 5 to get -3*-3+4 and got 14 that matches the 2nd number in (-3,14) so that is one of the answers.

OpenStudy (yourboydusty):

ok

OpenStudy (phi):

but we have to test the other choices does choice B work ? B. (3, -14) that says use 3 for x in -3 * x + 5 can you replace the x with 3? what do you get ?

OpenStudy (yourboydusty):

-4

OpenStudy (phi):

yes, and notice -4 is *not* -14 in (3,-14) so choice B is *not* an answer. can you check choice C ?

OpenStudy (phi):

to check C. (2, -1) what number do we use for x in -3*x + 5 ?

OpenStudy (yourboydusty):

-1 so yes

OpenStudy (phi):

-3*2 + 5 = -6+5 = -1 yes, choice C is also an answer one last choice to check. can you finish ?

OpenStudy (yourboydusty):

yes thank you

OpenStudy (yourboydusty):

and the last one is another answer

OpenStudy (phi):

yes, good job

OpenStudy (yourboydusty):

So its A C D

OpenStudy (3mar):

Great!